A garden hose having an internal diameter of 0.790 in. ( 2.0066 cm ) is connected to a lawn sprinkler that consists merely of an enclosure with 33 holes, each 0.055 in. ( 0.1397 cm ) in diameter. If the water in the hose has a speed of 3.67 ft/s ( 111.760 cm/s ), at what speed does it leave the sprinkler holes?
Answer #1

We know from the principle of continuity

A1 v1 = A2 v2

pi r1^2 v1^2 = pi r2^2 v2^2

r1^2 v1^2 = r2^2 v2^2

(2.0066/2)^2 x 111.760 = 30 (0.1397/2)^2 v2

v2 = (2.0066/2)^2 x 111.760/[30 (0.1397/2)^2] = 768.58 cm/s

Hence, v2 = 768.58 cm/s = 7.69 m/s = 25.22 ft/s